# Models, Methods and Waveforms for Estimation and Prediction of Doubly Sparse Time-Varying Channels

### v1

@@ Abstract
This paper investigates channel estimation for linear time-varying (LTV) wireless channels under double sparsity, i.e., sparsity in both the delay and Doppler domains. An on-grid approximation is first considered, enabling rigorous hierarchical-sparsity modeling and compressed sensing-based channel estimation. Guaranteed recovery conditions are provided for affine frequency division multiplexing (AFDM), orthogonal frequency division multiplexing (OFDM) and single-carrier modulation (SCM), highlighting the superiority of AFDM in terms of doubly sparse channel estimation. To address arbitrary Doppler shifts, a relaxed version of the on-grid model is introduced by making use of multiple elementary Expansion Models (BEM) each based on Discrete Prolate Spheroidal Sequences (DPSS). Next, theoretical guarantees are provided for the precision of this off-grid model before further extending it to tackle channel prediction by exploiting the inherent DPSS extrapolation capability. Finally, numerical results are provided to both validate the proposed off-grid model for channel estimation and prediction purposes under the double sparsity assumption and to compare the corresponding mean squared error (MSE) and the overhead performance when the different wireless waveforms are used.
Index Terms:
channel estimation, time-varying channels, AFDM, sparsity, basis expansion model, Slepian basis, DPSS, channel prediction
@@ I Introduction
Sparsity is an important feature of wireless channels, particularly in high-frequency bands. Sparsity emerges due to the limited number of dominant scatterers contributing to signal propagation and manifests both in the delay and the Doppler domains. By leveraging sparsity, advanced estimation methods can mitigate the challenges posed on channel estimation and prediction by high-mobility scenarios. The question thus arises as to 1) what delay-Doppler modeling can best leverage sparsity to enable efficient and low-complexity estimation and prediction of time-varying channels and 2) which wireless waveforms can make the best use of delay-Doppler sparsity e.g., for pilot overhead reduction.
@@
In wireless communication systems operating in sub-6GHz frequency bands, sparsity is often observed in the delay domain, where only a few dominant delay taps carry most of the channel energy [1]. Channel estimation methods leveraging this delay-domain sparsity have been proposed in various studies, such as [2] in which a grid-based discretization of the delay dimension is used to enable a compressive sensing framework. However, in higher-frequency bands, sparsity extends to the Doppler domain as well in scenarios such as vehicular or high-speed train communications. Delay-Doppler sparsity was assumed in [3] and leveraged to conceive enhanced channel estimation schemes for time-varying channels. However, sparsity was modeled as the sparsity of a one-dimensional array with no way to assign different sparsity levels to the delay and Doppler domains. The model in [4] also assumes a form of delay-Doppler sparsity, though under the restrictive assumption of one Doppler shift per delay tap. On-grid models, assuming that delay and Doppler shifts are quantized to a predefined grid, allow for structured sparse recovery methods. This quantization simplifies estimation but introduces grid mismatch errors [5]. To overcome the limitations of strict grid-based models, off-grid channel estimation techniques have been introduced by relaxing the assumption of discretized delay or Doppler shifts. Having estimates of the fractional part of Doppler frequency shifts enables straightforward channel extrapolation and, hence, prediction. Grid refinement techniques [6, 7] have been explored as a simple and low-complexity method to mitigate the performance degradation caused by grid mismatches. More advanced gridless approaches offering super-resolution recovery have also been investigated [8, 9, 3, 10]. Off-grid sparse Bayesian learning (SBL) [9, 3] is one such solution. Another solution based on atomic norm minimization is proposed in [10]. Such approaches can offer in principle better performance than grid refinement solutions but at the cost of higher computational complexity. Moreover, they cannot fully eliminate basis mismatch between real-world signals and the grid, leading to residual errors. They could also suffer from ill conditioning. This is particularly the case, as is argued in Sections IV and V, when the propagation link is characterized by a large number of “physical” channel paths contributing to each of the (refined) grid points. As an alternative, basis expansion models (BEM) can in principle be made to exploit sparsity. For example, the BEM based methods in [11, 12, 13] take advantage of sparsity, though only in the delay domain. There is thus a need to add support for Doppler domain sparsity to the BEM approach.
@@
Concerning waveforms conceived to deal with high-mobility scenarios, orthogonal time frequency space (OTFS) uses a two-dimensional (2D) modulation technique that transmits the information and pilot symbols in the delay-Doppler domain [14]. However, channel estimation overhead in OTFS cannot be reduced when the channel exhibits sparsity. Indeed, the performance gap between OTFS and OFDM narrows in favor of OFDM on channels with delay domain sparsity [14]. Another recently proposed waveform for communications in high-mobility scenarios is Affine Frequency Division Multiplexing (AFDM) [15]. While in AFDM data and pilot symbols are not directly transmitted in the delay-Doppler domain, AFDM can still reconstruct a delay-Doppler representation of the channel achieving full diversity on doubly dispersive channels. In the absence of sparsity, this leads to AFDM having a comparable bit error rate (BER) performance to that of OTFS but with the advantage of requiring less channel estimation overhead [15]. Under sparsity, our previous works [16, 17] proved the superiority of AFDM over both OFDM and OTFS in terms of pilot overhead using an on-grid channel model. In the current work, we build on those works and on a new sparsity modeling paradigm to expand our results to time-varying channel with off-grid Doppler frequency shifts and we include the study of channel prediction as a key application of this extension.
@@
    The statistical notion of delay-Doppler double sparsity (DS) for LTV channels is defined and rigorously linked to the hierarchical sparsity paradigm under a first on-grid approximation of Doppler shifts . This link is used to propose a sparse recovery algorithm for delay-Doppler profiles of wireless channels. Using hierarchical-sparsity mathematical tools, closed-form asymptotic results for the performance of this recovery is provided.
@@
    The effect of the waveform in use on DS-LTV channel estimation is analyzed using these theoretical results pointing to the superiority of AFDM over recovery schemes based on other waveforms in terms of channel estimation overhead.
@@
    A relaxation of the on-grid model to allow for off-grid Doppler frequency shifts is provided, resulting in a novel representation for LTV channels based on multiple elementary BEMs.
@@
    A new theoretical result is derived showing a logarithmic dependency in the inverse of the target mean-square error of the number of DPSS basis vectors needed to guarantee that target error when representing a samples vector from a narrow-band signal occupying a unit-frequency resolution bandwidth in the limit of large samples vector length. Such a signal is relevant because the channel component associated with a cluster of propagation paths sharing the same delay-Doppler grid point can modeled as one.
  5. 5.
    A prediction method based on extrapolating the basis vectors of the above BEM representation is finally proposed. Its optimality is established by proving it to be the reduced-rank MMSE estimator of a channel sample outside of the observation interval based on the knowledge of the channel delay-Doppler components within the interval.
@@
This paper is organized as follows. Section II provides a general mathematical framework for modeling doubly sparse time-varying wireless channels. Section III is dedicated to a first on-grid approximation of that model and to sparse-recovery channel estimation performance of different waveforms under that approximation. Section IV introduces a relaxation that allows for fractional Doppler frequency shifts along with a new model to capture the associated off-grid effects. Furthermore, the inherent properties of the proposed model are used to enable channel prediction. Finally, numerical results are given in Section V to validate the proposed models and the DS-LTV channel estimation and extrapolation performance of different waveform candidates.
@@
$\mathrm{Bernoulli}(p)$ is the Bernoulli distribution with probability $p$ and $\mathrm{B}(n,p)$ is the binomial distribution with parameters $(n,p)$. Notation $X\sim F$ means that random variable $X$ has distribution $F$. If $\mathcal{A}$ is a set, $|\mathcal{A}|$ stands for its cardinality. The set of all integers between $l$ and $m$ for some $(l,m)\in\mathbb{Z}^{2}$ (including $l$ and $m$) is denoted $\llbracket l\mkern 1.5mu..\mkern 1.5mum\rrbracket$. For a matrix $\mathbf{M}$, $[\mathbf{M}]_{c}$ stands for the $c$-th column. The ceiling operation is denoted as $\lceil$.$\rceil$ . ⌉. The modulo $N$ operation is denoted as $(\cdot)_{N}$. $\mathcal{F}$ stands for the continuous-time Fourier transform operator and $x\mapsto\mathrm{rect}(x)$ stands for the (standard) rectangular function. $\mathbf{I}_{N}$ is the $N\times N$ identity matrix.
II System Model: Linear time-varying channels
@@
$h(t,\tau)=\sum_{p=1}^{N_{\rm p}}g_{p}e^{\imath 2\pi\nu_{p}t\Delta f}\delta(\tau-\tau_{p}),$ (2)
is the continuous-time impulse response of the channel. Here, $N_{\rm p}\geq 1$ is the number of paths, $\delta(\cdot)$ is the Dirac delta function, and $g_{p},\nu_{p}$ and $\tau_{p}$ are the complex gain, Doppler shift (normalized with respect to the subcarrier spacing $\Delta f$), and the delay (normalized with respect to the sample period $T_{\rm s}$) associated with the $p$-th path, respectively. We define $\tau_{p}=l_{p}+\iota_{p}$ where $l_{p}\in\llbracket 0\mkern 1.5mu..\mkern 1.5muL-1\rrbracket$ is its integer part, while $\iota_{p}$ is the fractional part that satisfies $\frac{-1}{2}<\iota_{i}\leq\frac{1}{2}$. We also define $\nu_{p}=q_{p}+\kappa_{p}$ where $q_{p}\in\llbracket-Q\mkern 1.5mu..\mkern 1.5muQ\rrbracket$ is its integer part, while $\kappa_{p}$ is the fractional part satisfying $\frac{-1}{2}<\kappa_{i}\leq\frac{1}{2}$. In practice, the transmitted signal $s(t)$ is the continuous-time version of a discrete-time signal $s_{n}\triangleq s\left(nT_{\rm s}\right)$ generated from a vector $\mathbf{x}$ of $N$ symbols. These symbols could be either data symbols, pilot symbols, or a combination of both. Defining $r_{n}\triangleq r(nT_{\rm s})$ and $z_{n}\triangleq z(nT_{\rm s})$, we get the discrete-time version of the LTV channel model in (1)
$r_{n}=\sum_{p=1}^{N_{\rm p}}g_{p}e^{\imath 2\pi\nu_{p}n\Delta fT_{\rm s}}s(nT_{\rm s}-\tau_{p})+z_{n},\quad n\in\mathbb{Z}\>.$ (3)
@@
III Estimation of doubly sparse linear time-varying channels with the on-grid approximation
III-A First approximation: on-grid doubly sparse linear time-varying channels [16]
In a first approximation, we assume that both $\iota_{p}$ and $\kappa_{p}$ are zero and we define $L\triangleq\underset{p=1\cdots N_{\rm p}}{\max}\frac{\tau_{p}}{T_{\rm s}}+1$. Then the discrete-time LTV input-output model in (3) becomes
@@ III-A First approximation: on-grid doubly sparse linear time-varying channels [16]
The input-output relation in (4) defines an on-grid LTV channel with a $L-1$ maximum delay shift with the complex gain $h_{l,n}$ of the $l$-th path varying with time index $n$ as
$h_{l,n}=\sum_{p=1}^{N_{\rm p}}g_{p}e^{\imath 2\pi\frac{qn}{N}}\delta_{l-l_{p}},l=0\cdots L-1\>.$ (5)
@@
The number of paths $N_{\rm p}$, in (2) can now be expressed as $N_{\rm p}=\sum_{l=0}^{L-1}\sum_{q=-Q}^{Q}I_{l,q}$. Note that the distribution of the random variables $\left\{I_{l,q}\right\}_{l,q}$ controls the kind of sparsity the LTV channel might have. We opt for the distribution given by the following definition.
@@
The complex gain $h_{l,n}$ of the $l$-th path of the LTV channel varies with time as
@@
and there exist $10<p_{d},p_{D}<$ 1 such that
[…]
@@
Note that under Definition 1, $s_{\rm d}\triangleq\mathbb{E}\left[\sum_{l}I_{l}\right]=p_{d}L$ is the mean number of active delay taps in the delay-Doppler profile of the channel and can be thought of as the delay domain sparsity level while $s_{\rm D}\triangleq\mathbb{E}\left[\sum_{q}I_{q}^{(l)}\right]=p_{D}(2Q+1)$ is the mean number of active Doppler bins per delay tap and can be thought of as the Doppler domain sparsity level. Fig. 1 illustrates three different delay-Doppler sparsity models, fully described in [16] and dubbed Type-1, Type-2 and Type-3, that all fall under the scope of Definition 1 each with an additional different assumption on $I_{l}$ and $I_{q}^{(l)}$. Here, we just point out that the difference between Type-2 and Type-3 of Figures 1-(b) and 1-(c), respectively, is that in the latter the active Doppler bins per delay tap appear in clusters of random positions but of deterministic length as opposed to the absence of clusters in the former. The case where the delay taps have all the same (random) sparsity (as in Type-1 models of Fig. 1-(a)) also falls under Definition 1 by setting $I_{q}^{(l)}=I_{q}^{(0)},\forall l$.
@@
III-B Relation to hierarchical sparsity
Definition 2 (Hierarchical sparsity, [18]).
A vector $\mathbf{x}\in\mathbb{C}^{NM}$ is $\left(s_{N},s_{M}\right)$-sparse if it consists of $N$ blocks of size $M$, $s_{N}$ of which at most are non-zero, and each non-zero block is $s_{M}$-sparse.

### v2

# Models, Methods and Waveforms for Estimation and Prediction of Sparse Time-Varying Channels

@@ Abstract
This paper investigates channel estimation for linear time-varying (LTV) wireless channels under double sparsity, i.e., sparsity in both the delay and Doppler domains. An on-grid approximation is first considered, enabling rigorous hierarchical-sparsity modeling and compressed sensing-based channel estimation. Guaranteed recovery conditions are provided for affine frequency division multiplexing (AFDM), orthogonal frequency division multiplexing (OFDM) and single-carrier modulation (SCM), highlighting the superiority of AFDM in terms of doubly sparse channel estimation. To address arbitrary Doppler shifts, a relaxed version of the on-grid model is introduced by utilizing multiple elementary Expansion Models (BEM) each based on Discrete Prolate Spheroidal Sequences (DPSS). Next, theoretical guarantees are provided for the precision of this off-grid model before further extending it to tackle channel prediction by exploiting the inherent DPSS extrapolation capability. Finally, numerical results are provided to both validate the proposed off-grid model for channel estimation and prediction purposes under the double sparsity assumption and to compare the corresponding mean squared error (MSE) and overhead performance when different wireless waveforms are used.
@@ I Introduction
Sparsity is an important feature of wireless channels, particularly in high-frequency bands. Sparsity emerges due to the limited number of dominant scatterers contributing to signal propagation and manifests in both the delay and Doppler domains. By leveraging sparsity, advanced estimation methods can mitigate the challenges posed by high-mobility scenarios on channel estimation and prediction. The question thus arises as to 1) what delay-Doppler modeling can best leverage sparsity to enable efficient and low-complexity estimation and prediction of time-varying channels, and 2) which wireless waveforms can make the best use of delay-Doppler sparsity e.g., for pilot overhead reduction.
@@
In wireless communication systems operating in low-to-moderate mobility scenarios, sparsity is often observed in the delay domain, where only a few dominant delay taps carry most of the channel energy [1]. Sparsity extends to the Doppler domain as well in high-mobility scenarios such as vehicular or high-speed train communications. In both mobility regimes, sparsity is more accentuated in the beamspace domain of MIMO channels, especially when the number of antennas is large [2, 3, 4]. Channel estimation methods leveraging delay-domain sparsity have been proposed in various studies, such as [5] in which a grid-based discretization of the delay dimension is used to enable a compressive sensing framework. Delay-Doppler sparsity was assumed in [6] and leveraged to conceive enhanced channel estimation schemes for time-varying channels. However, sparsity was modeled as the sparsity of a one-dimensional array, with no way to assign different sparsity levels to the delay and Doppler domains. The model in [7] also assumes a form of delay-Doppler sparsity, though under the restrictive assumption of one Doppler shift per delay tap. On-grid models, assuming that delay and Doppler shifts are quantized to a predefined grid, allow for structured sparse recovery methods. This quantization simplifies estimation but introduces grid mismatch errors [8]. To overcome the limitations of strict grid-based models, off-grid channel estimation techniques have been introduced by relaxing the assumption of discretized delay or Doppler shifts. Having estimates of the fractional part of Doppler frequency shifts enables straightforward channel extrapolation and, hence, prediction. Grid refinement techniques [9, 10] have been explored as a simple and low-complexity method to mitigate the performance degradation caused by grid mismatches. More advanced gridless approaches offering super-resolution recovery have also been investigated [11, 12, 6, 13]. Off-grid sparse Bayesian learning (SBL) [12, 6] is one such solution. Another solution based on atomic norm minimization is proposed in [13]. Such approaches can, in principle, offer better performance than grid refinement solutions, but at the cost of higher computational complexity. Moreover, they cannot fully eliminate the basis mismatch between real-world signals and the grid, leading to residual errors. They could also suffer from ill conditioning. This is particularly the case, as argued in Section IV and validated in Section V, when the propagation link is characterized by a large number of “physical” channel paths contributing to each of the (refined) grid points. As an alternative, basis expansion models (BEM) do not attempt to estimate the individual off-grid delay-Doppler components but rather their combined effect, thus avoiding the ill-conditioning issue. Moreover, they can, in principle, be made to exploit sparsity. For example, the BEM based methods in [14, 15, 16] take advantage of sparsity, though only in the delay domain. There is thus a need to add support for Doppler domain sparsity to the BEM approach.
@@
Concerning waveforms conceived to deal with high-mobility scenarios, orthogonal time frequency space (OTFS) uses a two-dimensional (2D) modulation technique that transmits the information and pilot symbols in the delay-Doppler domain [17]. However, OTFS channel estimation overhead cannot be reduced when the channel exhibits sparsity [18]. Indeed, the performance gap between OTFS and OFDM narrows in favor of OFDM on channels with delay domain sparsity [19]. Another recently proposed waveform for communications in high-mobility scenarios is Affine Frequency Division Multiplexing (AFDM) [20]. While in AFDM, data and pilot symbols are not directly transmitted in the delay-Doppler domain; AFDM can still reconstruct a delay-Doppler representation of the channel to achieve full diversity on doubly dispersive channels. In the absence of sparsity, this leads to AFDM having a comparable bit error rate (BER) performance to that of OTFS, but with the advantage of requiring less channel estimation overhead [20]. Under sparsity, our previous works [21, 22] proved the superiority of AFDM over both OFDM and OTFS in terms of pilot overhead using an on-grid channel model. In the current work, we expand our results to time-varying channels with off-grid Doppler shifts and we address channel prediction as a key application of this extension.
@@
    The statistical notion of delay-Doppler double sparsity (DS) for LTV channels is defined and rigorously linked to the hierarchical sparsity paradigm under a first on-grid approximation of Doppler shifts . This link allows us to use hierarchical-sparsity mathematical tools from the compressed-sensing paradigm to analyze the effect of the used waveform on DS-LTV channel estimation, pointing to the superiority of AFDM over other waveforms in terms of sparse channel estimation overhead.
@@
    The on-grid model is relaxed to allow for off-grid Doppler shifts, resulting in a novel DS-LTV channel representation with theoretical precision guarantees based on multiple frequency-shifted elementary BEMs. This channel representation constitutes a novel approach to modeling delay-Doppler sparsity. In this approach, sparsity is not synonymous with relatively small numbers of channel paths but consists in assuming relatively small numbers of clusters of such paths, with each cluster being associated with a delay-Doppler grid point.
@@
    The proposed channel representation is then used to conceive a linear minimum mean square error (LMMSE) LTV channel estimator, replacing the more complex sparse-recovery approach that enabled us to obtain closed-form waveform performance analysis. The results of this theoretical waveform performance comparison, conducted under the on-grid approximation and the compressed-sensing paradigm, are shown using numerical results to continue to hold under the more realistic off-grid model and LMMSE channel estimation.
@@
    A prediction method based on extrapolating the basis vectors of the above BEM representation is finally proposed. Its optimality is established by linking it to the reduced-rank MMSE predictor.
@@
Section II provides a general mathematical framework for modeling doubly sparse, time-varying wireless channels. Section III is dedicated to a first on-grid approximation of that model and to the channel estimation performance of different waveforms under that approximation. Section IV introduces a relaxation that allows for fractional Doppler frequency shifts, along with a new model to capture the associated off-grid effects. Furthermore, the inherent properties of the proposed model are used to enable channel prediction. Finally, numerical results are provided in Section V to validate the proposed models and the DS-LTV channel estimation and extrapolation performance of different waveform candidates.
@@
$\mathrm{Bernoulli}(p)$ is the Bernoulli distribution with activation probability $p$ and $\mathrm{B}(n,p)$ is the binomial distribution with parameters $(n,p)$. Notation $X\sim F$ means that the random variable $X$ has a distribution $F$. If $\mathcal{A}$ is a set, $|\mathcal{A}|$ stands for its cardinality. The set of all integers between $l$ and $m$ for some $(l,m)\in\mathbb{Z}^{2}$ (including $l$ and $m$) is denoted $\llbracket l\mkern 1.5mu..\mkern 1.5mum\rrbracket$. For a matrix $\mathbf{M}$, $\mathbf{M}]_{c}$ stands for the $c$-th column. The ceiling operation is denoted as $\lceil.\rceil$. The modulo $N$ operation is denoted as $(\cdot)_{N}$. $\mathcal{F}$ stands for the continuous-time Fourier transform operator, and $x\mapsto\mathrm{rect}(x)$ stands for the (standard) rectangular function. $\mathbf{I}_{N}$ is the $N\times N$ identity matrix.
II System Model: Linear time-varying (LTV) channels
@@
$h(t,\tau)=\sum_{p=1}^{N_{\rm p}}g_{p}e^{\imath 2\pi\nu_{p}t\Delta f}\delta(\tau-\tau_{p}T_{\rm s}),$ (2)
is the continuous-time impulse response of the channel. Here, $N_{\rm p}\geq 1$ is the number of paths, $\delta(\cdot)$ is the Dirac delta function, $\Delta f$ is the subcarrier spacing, $T_{\rm s}$ is the sample period, $g_{p}$ is the complex gain of the $p$-th path, $\nu_{p}\triangleq\frac{\text{Doppler frequency in Hz}}{\Delta f}$ and $\tau_{p}\triangleq\frac{\text{Delay in seconds}}{T_{\rm s}}$. Moreover, we write $\tau_{p}=l_{p}+\iota_{p}$ with $l_{p}$ as the integer part of $\tau_{p}$, while $\iota_{p}$ is the fractional part that satisfies $\frac{-1}{2}<\iota_{i}\leq\frac{1}{2}$. Finally, $\nu_{p}=q_{p}+\kappa_{p}$, where $q_{p}\in\llbracket-Q\mkern 1.5mu..\mkern 1.5muQ\rrbracket$ is the integer part of $\nu_{q}$, while $\kappa_{p}$ is the fractional part satisfying $\frac{-1}{2}<\kappa_{i}\leq\frac{1}{2}$. In practice, the transmitted signal $s(t)$ is the continuous-time version of a discrete-time signal $s_{n}\triangleq s\left(nT_{\rm s}\right)$ generated from a vector $\mathbf{x}$ of $N$ symbols. These symbols could be either data symbols, pilot symbols, or a combination of both. Defining $r_{n}\triangleq r(nT_{\rm s})$ and $z_{n}\triangleq z(nT_{\rm s})$, we obtain the discrete-time version of the model in (1)
$r_{n}=\sum_{p=1}^{N_{\rm p}}g_{p}e^{\imath 2\pi\nu_{p}n\Delta fT_{\rm s}}s(nT_{\rm s}-\tau_{p}T_{\rm s})+z_{n},\quad n\in\mathbb{Z}\>.$ (3)
@@
III Estimation of doubly sparse LTV channels with the on-grid approximation
III-A First approximation: on-grid doubly sparse LTV channels
In a first approximation, we assume that both $\iota_{p}$ and $\kappa_{p}$ are zero, and we define $L\triangleq\underset{p=1\cdots N_{\rm p}}{\max}\tau_{p}+1$. The discrete-time input-output model in (3) becomes
@@ III-A First approximation: on-grid doubly sparse LTV channels
The input-output relation in (4) defines an on-grid LTV channel with a L1L-1 maximum delay shift with the complex gain $h_{l,n}$ of the $l$-th path varying with the time index $n$ as
$h_{l,n}=\sum_{p=1}^{N_{\rm p}}g_{p}e^{\imath 2\pi\frac{qn}{N}}\delta(l-l_{p}),l=0\cdots L-1\>.$ (5)
@@
The number of paths $N_{\rm p}$, in (2) can now be expressed as
$N_{\rm p}=\sum_{l=0}^{L-1}\sum_{q=-Q}^{Q}I_{l,q}\>.$ (8)
Note that the distribution of the random variables $\left\{I_{l,q}\right\}_{l,q}$ controls the type of sparsity that the LTV channel exhibits. We opt for the distribution given by the following definition.
@@
The complex gain $h_{l,n}$ of the $l$-th path varies with time as
@@
and there exist $0<p_{d},p_{D}<1$ such that
$I_{l,q}=I_{l}I_{q}^{(l)},\quad\forall(l,q)\in\llbracket 0\mkern 1.5mu..\mkern 1.5muL-1\rrbracket\times\llbracket-Q\mkern 1.5mu..\mkern 1.5muQ\rrbracket,$ (10)
@@
The use of the Bernoulli distribution to model channel sparsity is widely adopted [23]. It is a natural way to probabilistically model the activation or lack of activation of a channel tap, i.e., the presence or absence of an actual channel path at that tap position. Indeed, the sparsity level of a collection of channel taps can be directly linked to the activation probability of the Bernoulli distribution modeling each of them. More precisely, note that under Definition 1,
$s_{\rm d}\triangleq\mathbb{E}\left[\sum_{l}I_{l}\right]=p_{d}L$ (11)
is the mean number of active delay taps in the delay-Doppler profile of the channel and can be thought of as the delay domain sparsity level while
$s_{\rm D}\triangleq\mathbb{E}\left[\sum_{q}I_{q}^{(l)}\right]=p_{D}(2Q+1)$ (12)
is the mean number of active Doppler bins per delay tap and can be thought of as the Doppler domain sparsity level. Fig. 1 illustrates three different delay-Doppler sparsity models, fully described in [21] and dubbed Type-1, Type-2 and Type-3. All these models fall under the scope of Definition 1, each with an additional assumption on $I_{l}$ and $I_{q}^{(l)}$. For instance, in Type-3 models of Figure 1-(c), the active Doppler bins per delay tap appear in clusters of random positions but of deterministic length, as opposed to the absence of clusters in Type-2 models of Figure 1-(b). The case in which the delay taps have all the same (random) sparsity (as in Type-1 models of Fig. 1-(a)) also falls under Definition 1 by setting $I_{q}^{(l)}=I_{q}^{(0)},\forall l$.
@@
III-B Relation to hierarchical sparsity [24]
Definition 2 (Hierarchical sparsity).
Vector $\mathbf{x}\in\mathbb{C}^{NM}$ is $\left(s_{N},s_{M}\right)$-sparse if it consists of NN size-$M$ blocks, $s_{N}$ of which at most are non-zero, and each non-zero block is $s_{M}$-sparse.